OpenStudy (anonymous):

A computer byte is a string of 8 characters and each character is a 0 or a 1. An example of a computer byte is 01101011. Find the number of possible computer bytes.

OpenStudy (anonymous):

26

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

i just did that and got it correct

OpenStudy (anonymous):

the only options are 26, 56, 128 and 156

OpenStudy (anonymous):

ya its 26

OpenStudy (anonymous):

i mean 16 not 26

OpenStudy (anonymous):

26 isn't an option

OpenStudy (anonymous):

that is what i meant my bad i know it was the first response

OpenStudy (anonymous):

its really easy

OpenStudy (anonymous):

I just have to be sure, but thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

use the sequence 8x+1-98+78-y+mx=b+x-y=x2